# High-contrast double Bragg interferometry via detuning control

### v1

@@ Abstract
We propose high-contrast Mach-Zehnder atom interferometers based on double Bragg diffraction (DBD) operating under external acceleration. To mitigate differential Doppler shifts and experimental imperfections, we introduce a tri-frequency laser scheme with dynamic detuning control. We evaluate four detuning-control strategies—conventional DBD, constant detuning, linear detuning sweep (DS-DBD), and a hybrid protocol combining detuning sweep with optimal control theory (OCT)—using exact numerical simulations and a five-level S-matrix model. The OCT strategy provides the highest robustness, maintaining contrast above 95% under realistic conditions, while the DS-DBD strategy sustains contrast above 90% for well-collimated Bose-Einstein condensates. These results offer practical pathways to enhancing DBD-based interferometers for precision quantum sensing and fundamental physics tests.
@@ I Introduction
Atom interferometry (AI) enables precision measurements of inertial and fundamental physical quantities by coherently splitting and recombining atomic wave packets along distinct paths. Applications include atomic gravimetry Peters, Chung, and Chu (1999, 2001); Hu et al. (2013); Altin et al. (2013); Poli et al. (2011); Szigeti et al. (2020); Snadden et al. (1998); McGuirk et al. (2002); Rosi et al. (2015), rotation and inertial sensing Gustavson, Bouyer, and Kasevich (1997); Stockton, Takase, and Kasevich (2011); Geiger et al. (2011); Gautier et al. (2022); d’Armagnac de Castanet et al. (2024); Stolzenberg et al. (2025); Pelluet et al. (2025), precision measurements of the fundamental constants Fixler et al. (2007); Lamporesi et al. (2008); Rosi et al. (2014); Parker et al. (2018); Morel et al. (2020) and searches for new physics beyond the Standard Model such as the detection of ultralight dark matter Geraci and Derevianko (2016); Arvanitaki et al. (2018); Stadnik and Flambaum (2016); Graham et al. (2016); Du et al. (2022). The precision of these measurements hinges on extracting the phase accumulated by atoms in quantum superposition—either of motional or internal states—via fitting the signal to an a priori theoretical model.
Double Bragg diffraction (DBD) is a large-momentum-transfer technique used in atom interferometry that couples atoms to symmetric momentum states via two counter-propagating optical lattices with orthogonal polarizations (Fig. 1, left panel), first demonstrated by the group of Rasel and his collaborators within the QUANTUS consortium Ahlers et al. (2016). Compared to single Bragg diffraction (SBD) and Raman schemes, DBD offers a doubled interferometric scale factor at a given order while operating entirely within a single internal state, thereby avoiding the additional decoherence channels inherent to Raman transitions Ozeri et al. (2005); Uys et al. (2010). Moreover, its intrinsic parity symmetry Li et al. (2024) suppresses laser phase noise and systematic effects, making it well-suited for microgravity applications Barrett et al. (2016); Becker et al. (2018); Gaaloul et al. (2022); Elliott et al. (2023) and horizontal geometries Gebbe et al. (2021); Ahlers et al. (2016). However, the multi-level nature of Bragg transitions introduces unwanted parasitic paths and diffraction phases Hartmann et al. (2020); Kirsten-Siemß et al. (2023), reducing the efficiencies of the beam-splitter (BS) and the mirror (M) operations and thus limiting the overall contrast of DBD-based interferometers Ahlers et al. (2016); Gebbe et al. (2021); Stolzenberg et al. (2025).
@@
Figure 1: Schematics of a double Bragg atom interferometer under microgravity. Left: Experimental setup of a DBD pulse using counter-propagating optical lattices $L_{1}$ and $L_{2}$ with orthogonal polarizations $\hat{\sigma}_{1}$ and $\hat{\sigma}_{2}$. Right: Real-space atomic density evolution $|\psi(z,t)|^{2}$, normalized to the initial maximum $|\psi_{max}|^{2}=\max_{z}|\psi(z,0)|^{2}$ and shown in decibel units, for conventional (C-DBD) and optimized (OCT) Mach-Zehnder interferometers with phase shifts $\Delta\phi=0$ (left column) and $\pi$ (right column), adjusted via the interrogation time $T$. Atomic density are extracted from exact numerical simulations. Red shaded regions mark the three DBD pulses.
@@
Interferometric contrast (or visibility) is a key factor determining the sensitivity and operational limits of atom interferometers. It directly affects the maximum achievable interrogation time $T$ and effective momentum transfer $k_{\text{eff}}$, beyond which contrast loss exceeds the measurement uncertainty. At the shot-noise limit, the acceleration sensitivity of a Mach-Zehnder (MZ) interferometer scales as $\delta a=1/(\sqrt{N}k_{\text{eff}}T^{2})$, where $N$ is the number of uncorrelated atoms Bordé (1989); Kasevich and Chu (1991); Torii et al. (2000); Kasevich and Chu (1992); Storey and Cohen-Tannoudji (1994); Schleich, Greenberger, and Rasel (2013); Bongs, Launay, and Kasevich (2006); Kritsotakis et al. (2018). While near-unity contrast can be achieved in SBD-based interferometers Torii et al. (2000), DBD-based schemes are typically more sensitive to polarization errors and momentum spread, which degrade pulse efficiency and overall interferometric contrast.
In this work, we theoretically study double Bragg diffraction atom interferometers operating under a constant external acceleration, such as gravity. In Sec. II, we propose a tri-frequency laser configuration with time-dependent detuning to compensate the differential Doppler shift and improve pulse efficiency. Four detuning-control strategies are introduced: conventional DBD, constant detuning, a linear detuning sweep, and a hybrid protocol that incorporates optimal control theory (OCT) Peirce, Dahleh, and Rabitz (1988); Riahi et al. (2016); Kovachy, Chiow, and Kasevich (2012); Louie et al. (2023); Saywell et al. (2023); Goerz, Kasevich, and Malinovsky (2023); Rodzinka et al. (2024). In Sec. III, we compare the robustness of the four strategies under realistic experimental conditions. With the OCT strategy, we successfully eliminate most parasitic populations and recover near-unity contrast compared to the conventional DBD scheme (see Fig. 1, right panel). Our results offer practical means for enhancing contrast and robustness in DBD-based interferometers, with direct relevance to precision quantum sensing applications in both ground and space-based platforms.
@@ II.1 Tri-frequency double Bragg accelerometers
To measure a constant linear acceleration $\bm{g}=g\hat{z}$ along the $z$-axis, we consider a MZ atom interferometer based on DBD. In microgravity or weak-acceleration regimes, a standard dual-frequency scheme suffices to drive symmetric momentum transfer in both directions Gebbe et al. (2021); Ahlers et al. (2016). However, under strong acceleration—such as gravity—the resulting differential Doppler shift $\nu_{g}=2\,k_{L}gt$ (where $k_{L}$ is the laser wave number) breaks this symmetry, preventing simultaneous resonance with both upward and downward Bragg transitions. To overcome this limitation, we propose a tri-frequency retro-reflective setup (see Fig. 2(a)), previously demonstrated in double Raman gravimeters Malossi et al. (2010). In this configuration, one of the two input frequencies (e.g., $\omega_{b}$) is replaced by a pair of frequencies $\omega_{b}\pm\nu_{D}$, where the detuning $\nu_{D}=2k_{L}a_{L}t$ linearly compensates the Doppler shift $\nu_{g}$. Due to momentum selectivity, the accelerating atoms selectively couple to four of the six laser beams, forming a pair of resonant Bragg lattices, as highlighted by the red box in Fig. 2(a).
The one-dimensional (1D) single-particle Hamiltonian describing the tri-frequency double Bragg diffraction of an atom with mass $m$ moving under a strong constant acceleration $g$, as illustrated in Fig. 2(a-b), can be written in the laboratory frame as
@@
where $g$ points in the positive $z$-direction, and $\varepsilon_{pol}=|\sigma_{\perp}\cdot\sigma_{\parallel}|$ quantifies the polarization error arising from imperfect beam polarizations. $\Omega(t)$ is an effective two-photon Rabi frequency for the light-atom interaction, which can be time-dependent. For a MZ atom interferometer with triple Gaussian pulses (see Fig. 1, right panel), $\Omega(t)$ takes the following form:
@@
Figure 2: (a) Tri-frequency laser configuration enabling double Bragg diffraction of atoms under constant acceleration $\bm{g}=g\hat{z}$, with dynamic Doppler shift compensation. The red box highlights four beams that resonantly drive DBD as the Doppler shift increases. (b) Energy level diagram of tri-frequency DBD including Doppler and AC-Stark shifts. The upward and downward Bragg transitions are driven by frequency differences $\Delta\omega_{1,2}=\omega_{b}-\omega_{a}\pm\nu_{D}$, where $\nu_{D}$ is dynamically tuned to compensate the Doppler shift $\nu_{g}=2k_{L}gt$.
It should be noted that $\Delta\omega(t)t=4\,\omega_{\mathrm{rec}}t+\Delta(t)t\equiv\phi(t_{i})+\int_{t_{i}}^{t}[\omega_{b}(t)-\omega_{a}(t)]\,dt$ in Eq. (S1) (with $\omega_{\mathrm{rec}}\equiv\hbar k_{L}^{2}/(2m)$) denotes the physically accumulated phase difference between the blue (or purple) and red lasers shown in Fig. 2(a-b), evaluated in the COM frame of the twin Bragg lattices where $\nu_{D}=0$ (see footnote [44] in Ref. Li et al. (2024)). In this COM frame, the DBD Hamiltonian (S1) transforms to
@@
via a time-dependent unitary transformation $H_{COM}(t)=UH_{lab}(t)U^{\dagger}+i\hbar\dot{U}U^{\dagger}$ (see supplementary material for details). $a_{L}=\dot{\nu}_{D}/(2k_{L})$ is the COM acceleration of the twin Bragg lattices which can be fine-tuned to approach the linear acceleration $g$ such that $g-a_{L}$ is approaching zero. From Eq. (3), one can identify an effective acceleration in the new frame as $g_{eff}\equiv g-a_{L}$. For the remainder of this paper, we denote the effective acceleration $g_{eff}$ simply as $g$, and always work in the twin-lattice COM frame unless otherwise specified.
Under the microgravity assumption in the accelerated frame, i.e., $|mgz|\ll 2\hbar\Omega_{R}$ and $|g|\ll\hbar k_{L}/(mt)$ for all relevant times $t$ and positions $z$ during the interferometer sequence, the last term in Hamiltonian (3) can be treated as a perturbation and neglected during the light-matter interaction. This allows the application of the double Bragg theory developed in Ref. Li et al. (2024). During free evolution intervals between pulses, the effective acceleration leads to both a momentum shift of the atomic wave packet and an additional propagation phase dependent on the initial momentum state $|p\rangle$ and the free evolution time $T$:
@@ II.2.1 Conventional-DBD Mach-Zehnder protocol
Figure 3: Comparison of beam-splitter (BS) and mirror (M) efficiencies for four strategies: conventional DBD (C-DBD), constant-detuning DBD (CD-DBD), linear-detuning-sweep DBD (DS-DBD), and a hybrid strategy combining DS-BS with OCT-mirror (OCT). Efficiencies are evaluated for an input Gaussian wave packet with a momentum width $\sigma_{p}=0.05\,\hbar k_{L}$, centered at $p_{0}=0$ for BS and $p_{0}=2\,\hbar k_{L}$ for M.
@@ II.2.3 Linear-detuning-sweep mitigated Mach-Zehnder protocol
This approach addresses the momentum-dependent energy shift arising from time-varying AC-Stark shifts, Doppler effects due to finite $\sigma_{p}$ or nonzero $p_{0}$, and polarization errors $\varepsilon_{pol}$—all of which may fluctuate shot-to-shot and render the resonance condition analytically intractable. To circumvent the need for precise resonance knowledge, we adopt a detuning control strategy inspired by the principle of adiabatic passage in two-level systems Steck (2025); Landau (1932); Zener and Fowler (1932), a well-established technique in nuclear magnetic resonance Bloch and Siegert (1940); Bloch (1946); Abragam (1961); Baum, Tycko, and Pines (1985); Philpott (1987). We extend this concept to the multi-level case of double Bragg diffraction via a Magnus expansion in the quasi-Bragg regime Li et al. (2024), yielding an effective two-level model between $|0\rangle=|0\hbar k_{L}\rangle$ and $|1\rangle=(|2\hbar k_{L}\rangle+|-2\hbar k_{L}\rangle)/\sqrt{2}$ (see Fig. 4(a)).
@@
A linear sweep then enables robust and highly efficient population transfer (illustrated in Fig. 4(b)), achieving $\eta_{BS}=99.937\%$ for a DS beam-splitter at $\varepsilon_{pol}=0$—a substantial improvement over the C-DBD and within one per mil of unity, outperforming CD-DBD (Fig. 3). In contrast, the DS mirror offers only a modest gain of about $1\%$, with an efficiency $\eta_{M}=97.465\%$, still well below unity. Alternative monotonic detuning profiles (e.g., sigmoid) were also explored but offered no significant gain. The resulting DS-DBD interferometer follows a BS–M–BS sequence with interrogation time $T$. This approach is effective when the range of polarization errors is known (e.g., $\varepsilon_{pol}<5\%$), and the uncertainties in $p_{0}$ and $\sigma_{p}$ are small and under control (e.g., for well-collimated BECs).

### v2

@@ Abstract
We propose high-contrast Mach-Zehnder atom interferometers based on double Bragg diffraction (DBD) operating under external acceleration. To mitigate differential Doppler shifts and experimental imperfections, we introduce a tri-frequency laser scheme with dynamic detuning control. We evaluate four detuning-control strategies—conventional DBD, constant detuning, linear detuning sweep (DS-DBD), and a hybrid protocol combining detuning sweep with optimal control theory (OCT)—using exact numerical simulations and a five-level S-matrix model. The OCT strategy provides the highest robustness, maintaining contrast above 95% under realistic conditions, while the DS-DBD strategy sustains contrast above 90% for well-collimated Bose-Einstein condensates. These results offer practical pathways to high-contrast, large-momentum-transfer DBD-based interferometers for precision quantum sensing and fundamental physics tests.
@@ I Introduction
Atom interferometry (AI) enables precision measurements of inertial and fundamental physical quantities by coherently splitting and recombining atomic wave packets along distinct paths. Applications range from atomic gravimetry Peters et al. (1999, 2001); Poli et al. (2011); Hu et al. (2013); Altin et al. (2013); Karcher et al. (2018); Szigeti et al. (2020) and gravity gradiometry Snadden et al. (1998); McGuirk et al. (2002); Rosi et al. (2015), rotation and inertial sensing Gustavson et al. (1997); Stockton et al. (2011); Geiger et al. (2011); Gautier et al. (2022); d’Armagnac de Castanet et al. (2024); Stolzenberg et al. (2025); Pelluet et al. (2025), to precision determinations of the fundamental constants Fixler et al. (2007); Lamporesi et al. (2008); Rosi et al. (2014); Parker et al. (2018); Morel et al. (2020) and searches for new physics beyond the Standard Model such as ultralight dark matter Geraci and Derevianko (2016); Arvanitaki et al. (2018); Stadnik and Flambaum (2016); Graham et al. (2016); Du et al. (2022). A central challenge in advancing interferometer performance is achieving large momentum transfer (LMT) while preserving high contrast.
Double Bragg diffraction (DBD) is an LMT technique used in atom interferometry that couples atoms to symmetric momentum states via two counter-propagating optical lattices with orthogonal polarizations (Fig. 1, left panel), first demonstrated by the group of Ernst Maria Rasel and his collaborators within the QUANTUS consortium Ahlers et al. (2016). Compared to single Bragg and Raman schemes Torii et al. (2000); Bordé-PLA-1989; Kasevich and Chu (1991, 1992), DBD doubles the interferometric scale factor at a given order while operating entirely within one internal state, eliminating hyperfine-state decoherence channels Ozeri et al. (2005); Uys et al. (2010). Relative to double Raman diffraction (DRD) Lévèque et al. (2009), DBD is significantly simpler to implement experimentally: its required two-photon detuning typically lies in the kHz regime (rather than GHz), unwanted acousto-optic-modulator-generated sidebands are easily removed in free space, and the same laser system can drive subsequent Bloch oscillations (BO) for momentum transfer at the $10^{2}$$10^{3}\,\hbar k_{L}$ scale Gebbe et al. (2021), where $k_{L}$ is the laser wave number. Moreover, higher-order single diffraction up to $24\,\hbar k_{L}$ Müller et al. (2008) and sequential diffractions up to $102\,\hbar k_{L}$ Chiow et al. (2011) have only been demonstrated with Bragg diffraction to date. DBD also exhibits an intrinsic parity symmetry Li et al. (2024) that suppresses common-mode phase noise, making it particularly suitable for horizontal geometries and microgravity platforms Barrett et al. (2016); Becker et al. (2018); Gaaloul et al. (2022); Elliott et al. (2023). However, the conventional retro-reflected DBD configuration Ahlers et al. (2016); Gebbe et al. (2021) cannot operate in non-degenerate settings (e.g., vertical geometries under gravity), which limits its competitiveness and practical applicability.
@@
Figure 1: Schematics of a double Bragg atom interferometer under microgravity. Left: Experimental setup of a DBD pulse using counter-propagating optical lattices $L_{1}$ (red shaded) and $L_{2}$ (blue shaded) with orthogonal polarizations $\hat{\sigma}_{1}$ and $\hat{\sigma}_{2}$. Right: Atomic density evolution in the twin-lattice center-of-mass frame $|\psi(z,t)|^{2}$, normalized to its initial maximum $|\psi_{max}|^{2}=\max_{z}|\psi(z,0)|^{2}$ and shown in decibel units, for conventional (C-DBD) and optimized (OCT) Mach-Zehnder interferometers with phase shifts $\Delta\phi=0$ (left column) and $\pi$ (right column), adjusted via the interrogation time $T$. Atomic densities are obtained from exact numerical simulations, with red-shaded regions indicating the three DBD pulses. Only the visibly populated diffracted ports are labeled in the right panel after the last DBD pulse.
@@
Interferometric contrast (or visibility) is a central figure of merit determining both the sensitivity and operational limits of atom interferometers. It determines the maximum usable interrogation time $T$ and effective momentum transfer $k_{\text{eff}}$, beyond which contrast loss dominates the measurement uncertainty. For a Mach–Zehnder (MZ) atom interferometer operating at the shot-noise limit, the acceleration sensitivity is given by $\delta a=1/(\mathbf{C}\sqrt{N}\,k_{\mathrm{eff}}T^{2})$, where $N$ is the number of uncorrelated atoms detected and $\mathbf{C}$ is the contrast (or visibility) Storey and Cohen-Tannoudji (1994); Schleich et al. (2013); Hardman et al. (2014); Kritsotakis et al. (2018). While near-unity contrast can be achieved in SBD-based interferometers Torii et al. (2000), DBD-based schemes are typically more susceptible to polarization errors, momentum spreads, and parasitic-path phase shifts Hartmann et al. (2020); Kirsten-Siemß et al. (2023); Lahuerta et al. (2025), which reduce the efficiency of both the beam-splitter (BS) and—more severely—the mirror (M) pulses, ultimately limiting the achievable contrast in state-of-the-art DBD interferometers Ahlers et al. (2016); Gebbe et al. (2021); Stolzenberg et al. (2025).
In this work, we extended double Bragg diffraction to regimes involving large external accelerations—relevant for applications such as gravimetry—by introducing a third, dynamically tunable laser frequency that restores symmetry between the upward and downward Bragg transitions in Sec. II. We develop a microscopic description of the full DBD Mach–Zehnder interferometer using a five-level S-matrix formalism and compare the robustness of different detuning-control strategies in Sec. III. Building upon the detuning-sweep (or adiabatic passage) and optimal-control-theory (OCT) methods introduced in the previous works Kovachy et al. (2012); Li et al. (2024); Peirce et al. (1988); Riahi et al. (2016); Saywell et al. (2018, 2020); Louie et al. (2023); Saywell et al. (2023); Goerz et al. (2023); Rodzinka et al. (2024); LeDesma et al. (2024), we systematically generalize these techniques to include the DBD mirror pulse—historically the main performance bottleneck. We find that OCT yields a mirror pulse achieving more than $99.5\%$ efficiency for an ultracold atomic ensemble with momentum width $0.05\,\hbar k_{L}$. Together, these developments provide a realistic and experimentally feasible route toward robust high-contrast DBD interferometers, with the potential to exceed $95\%$ overall contrast under realistic imperfections using advanced detuning-controlled pulses, expanding the applications of double Bragg diffraction in precision quantum sensing on both terrestrial and space-based platforms.
@@ II.1 Tri-frequency double Bragg accelerometers
To measure a constant external acceleration $\bm{g}=g\hat{z}$, we consider a Mach–Zehnder atom interferometer based on DBD. This constant-acceleration assumption is fully compatible with real-world applications—such as gravimetry, geodesy, and inertial navigation—where the acceleration may vary between shots but is effectively constant over a single interferometer cycle. In microgravity or weak-acceleration regimes, a standard dual-frequency scheme suffices to drive symmetric momentum transfer in opposite directions Gebbe et al. (2021); Ahlers et al. (2016). However, under strong acceleration—such as gravity—the resulting differential Doppler shift $\nu_{g}=2\,k_{L}gt$ (frequencies are given in angular units unless stated otherwise) breaks this symmetry, preventing simultaneous resonance with the upward and downward Bragg transitions. To overcome this limitation, we propose a tri-frequency retro-reflective configuration (see Fig. 2(a)), previously demonstrated in double Raman gravimeters Malossi et al. (2010). In this configuration, one of the two input frequencies (e.g., $\omega_{b}$) is replaced by a frequency pair $\omega_{b}\pm\nu_{D}$, where the detuning $\nu_{D}=2k_{L}a_{L}t$ linearly compensates the Doppler shift $\nu_{g}$. Due to momentum selectivity Li et al. (2024), the accelerating atoms couple only to four of the six beams, forming a pair of resonant Bragg lattices highlighted in Fig. 2(a).
The one-dimensional (1D) single-particle Hamiltonian describing tri-frequency double Bragg diffraction of an atom with mass $m$ subjected to a strong constant acceleration $\bm{g}=g\hat{z}$, as illustrated in Fig. 2(a–b), can be written in the laboratory frame as
@@
where $g>0$ and $\varepsilon_{pol}=|\mathbf{\sigma_{\perp}}^{\dagger}\mathbf{\sigma_{\parallel}}|$ quantifies the polarization error arising from imperfect optical-field polarizations. For a conservative (worst-case) estimate of polarization-error effects, we neglect the temporal modulation associated with this term. A detailed derivation of Eq. (1) is provided in the Supplementary Material. $\Omega(t)$ is an effective two-photon Rabi frequency for the light-atom interaction, which can be time-dependent. For an MZ atom interferometer with triple Gaussian pulses (see Fig. 1, right panel), $\Omega(t)$ takes the following form:
@@
Figure 2: (a) Tri-frequency laser configuration enabling double Bragg diffraction of atoms under a strong constant acceleration $\bm{g}=g\hat{z}$, with dynamic Doppler shift compensation. The red box highlights four beams that resonantly drive DBD as the Doppler shift increases. (b) Corresponding energy-level diagram including Doppler and AC-Stark shifts. Upward and downward Bragg transitions are driven by frequency differences $\Delta\omega_{1,2}=\omega_{b}-\omega_{a}\pm\nu_{D}$, where $\nu_{D}$ is dynamically tuned to compensate the Doppler shift $\nu_{g}=2k_{L}gt$. In (b), the solid lines indicate resonant transitions, whereas dashed lines represent off-resonant couplings.
It should be noted that $\Delta\omega(t)t=4\,\omega_{\mathrm{rec}}t+\Delta(t)t\equiv\phi(t_{i})+\int_{t_{i}}^{t}[\omega_{b}(t)-\omega_{a}(t)]\,dt$ in Eq. (1), with $\omega_{\mathrm{rec}}\equiv\hbar k_{L}^{2}/(2m)$, denotes the physically accumulated phase difference between the blue (or purple) and red lasers shown in Fig. 2(a-b), evaluated in the center-of-mass (COM) frame of the twin Bragg lattices where $\nu_{D}=0$. In this COM frame, the Hamiltonian in Eq. (1) transforms into
@@
via a time-dependent unitary transformation $H_{COM}(t)=UH_{lab}(t)U^{\dagger}+i\hbar\dot{U}U^{\dagger}$ (see Supplementary Material for details). $a_{L}=\dot{\nu}_{D}/(2k_{L})$ is the COM acceleration of the twin Bragg lattices which can be fine-tuned to approach the linear acceleration $g$ such that $g-a_{L}$ is approaching zero. From Eq. (3), one identifies an effective acceleration in the transformed frame, $g_{eff}\equiv g-a_{L}$. For the remainder of this paper, we refer to $g_{eff}$ simply as $g$, and perform all calculations in the twin-lattice COM frame unless otherwise specified.
Under the microgravity assumption in the transformed frame, i.e., $|mgz|\ll 2\hbar\Omega_{R}$ and $|g|\ll\hbar k_{L}/(mt)$ for all relevant times $t$ and positions $z$ during the interferometer sequence, the last term in Hamiltonian (3) can be treated as a perturbation and neglected during the light-matter interaction. This allows the application of the double Bragg theory developed in Ref. Li et al. (2024). During free evolution intervals between pulses, the effective acceleration leads to both a momentum shift of the atomic wave packet and an additional propagation phase dependent on the initial momentum state $|p\rangle$ and the free evolution time $T$:
@@ II.2.1 Conventional-DBD Mach-Zehnder protocol
Figure 3: Comparison of beam-splitter (BS) and mirror (M) inefficiencies ($1-\eta_{BS|M}$) for four strategies: conventional DBD (C-DBD), constant-detuning DBD (CD-DBD), linear-detuning-sweep DBD (DS-DBD), and a hybrid strategy combining DS-BS with OCT-M (OCT). BS and M efficiencies $\eta_{BS|M}$ are evaluated for an input Gaussian wave packet with a momentum width $\sigma_{p}=0.05\,\hbar k_{L}$, centered at $p_{0}=0$ for BS and $p_{0}=2\,\hbar k_{L}$ for M without polarization error ($\varepsilon_{pol}=0$).
@@ II.2.3 Linear-detuning-sweep mitigated Mach-Zehnder protocol
This approach addresses the momentum-dependent energy shift arising from time-varying AC-Stark shifts, Doppler effects due to finite $\sigma_{p}$ or nonzero $p_{0}$, and polarization errors $\varepsilon_{pol}$—all of which may fluctuate shot-to-shot and render the resonance condition analytically intractable. To circumvent the need for precise resonance knowledge, we adopt a detuning control strategy inspired by the principle of adiabatic passage in two-level systems Steck (revision 0.16.4, 7 May 2025); Landau (1932); Zener and Fowler (1932), a well-established technique in nuclear magnetic resonance Bloch and Siegert (1940); Bloch (1946); Abragam (1961); Baum et al. (1985); Philpott (1987). We extend this concept to the multi-level case of double Bragg diffraction via a Magnus expansion in the quasi-Bragg regime Li et al. (2024), yielding an effective two-level model between $|0\rangle=|0\hbar k_{L}\rangle$ and $|1\rangle=(|2\hbar k_{L}\rangle+|-2\hbar k_{L}\rangle)/\sqrt{2}$ (see Fig. 4(a)).
@@
A linear detuning sweep enables robust and highly efficient population transfer (illustrated in Fig. 4(b)), yielding a beam-splitter efficiency of $\eta_{BS}=99.937\%$ at $\varepsilon_{pol}=0$. This represents a substantial improvement over both C-DBD and CD-DBD, with a residual inefficiency below $10^{-3}$ (see Fig. 3). In contrast, the DS mirror offers only a modest gain of about $1\%$, achieving an efficiency of $\eta_{M}=97.465\%$. This indicates that substantial room for further optimization via OCT remains. Alternative monotonic detuning profiles (e.g., sigmoid) were also explored but offered no significant gain. The resulting DS-DBD interferometer follows a BS–M–BS sequence with interrogation time $T$. This approach is effective when the range of polarization errors is known (e.g., $\varepsilon_{pol}<5\%$), and the uncertainties in $p_{0}$ and $\sigma_{p}$ are small and under control (e.g., for well-collimated BECs).